On a part-time job, you are asked to bring a cylindrical iron rod of density 7800 kg/m3 , length 89.5 cm and diameter 2.40 cm from a storage room to a machinist. Calculate the weight of the rod, w. The acceleration due to gravity, g = 9.81 m/s
Answer #1

Area of rod = pi*(0.024/2)^2*0.895

mass = area * density

= pi*(0.024/2)^2*0.895*7800

Weight = mass*gravity

= pi*(0.024/2)^2*0.895*7800*9.81

= 30.98 N